# Greedy Heuristics for Sampling-based Motion Planning in High-Dimensional State Spaces

### v1

@@ 7 Conclusion and Discussion
Anytime sampling-based planning algorithms have been proven to perform better in high-dimensional state spaces compared to graph-search approaches since these methods do not explicitly construct high-dimensional occupancy obstacles; instead, they rely on random sampling to asymptotically find optimal solutions. Algorithms like RRT* increasingly build dense approximations over the problem domain based on randomly sampled states and find continually improving solutions over time. This makes RRT* probabilistically complete and almost-surely asymptotically optimal (Karaman et al., 2011).
@@
Although they are effective in continuous planning problems, their search is often inefficient since they tend to uniformly explore the entire state space (i.e., they are space-filling algorithms). In other words, these algorithms tend to spend time exploring unnecessary regions of the problem domain where solutions do not exist. This wasted computational effort becomes more pronounced in planning problems with higher dimensions of state space.
@@
Previous works therefore attempt to focus the search solely on promising regions of the problem domain where solutions are likely to exist. Algorithms such as Informed RRT* directly focus sampling on a subset of the state space determined by the current solution path. This $L^{2}$ informed set contains the set of states that can potentially provide better solutions and often ensures faster convergence guarantees for planning scenarios aimed at minimizing path lengths. However, since this set is estimated using the cost of the current solution path, it can still result in larger sampling regions, especially when the current solution contains many twists and turns.
@@
This paper presents an alternative version of the informed set, called the $L^{2}$ greedy informed set, which is approximated based on the exploited knowledge of the current solution path using the maximum admissible heuristic cost (Section 3). Instead of using the cost of the current solution as an admissible heuristic for the informed set, the greedy informed set only considers the maximum heuristic cost within the current solution path. This results in smaller informed subsets and is shown to have better convergence properties than the informed set, typically in problems with complex initial solution paths.
@@
While direct informed sampling techniques prove effective compared to traditional space-filling approaches, their effectiveness only manifests once an initial solution is found. Consequently, the benefits of focused sampling become irrelevant in complex planning problems where finding an initial solution is challenging, especially in high-dimensional state spaces. Hence, finding a fast initial solution is crucial for sampling-based planning algorithms to leverage all the benefits of informed sets. This paper also presents an algorithm that leverages bi-directional search to accelerate the process of finding initial solutions.
@@
To this end, this paper presents an almost-surely asymptotically optimal path planning algorithm that utilizes greedy heuristic techniques to address challenges in high-dimensional planning domains (Section 4). G-RRT* maintains two rapidly growing random geometric graphs (RGGs) to explore the problem state space and uses a greedy connection heuristic to guide the trees toward each other. This helps G-RRT* find faster initial solutions and often escape local minima with biased Voronoi regions.
@@
G-RRT* also focuses the search on the $L^{2}$ greedy informed set once the initial solution is found, i.e., forward and reverse search trees meet each other. This allows G-RRT* to search efficiently on the smaller subset of the problem domain with exploited knowledge the current solution path. Moreover, a greedy biasing ratio is introduced to balance uniform sampling from the problem domain (i.e., exploration with $L^{2}$ informed set) for optimal guarantees and path-biased sampling (i.e., exploitation with $L^{2}$ greedy informed set) for accelerating convergence. As a result, G-RRT* is probabilistically complete and almost-surely asymptotically optimal (Section 5).
@@
The benefits of the greedy heuristic techniques are demonstrated on abstract planning problems, manipulation problems in simulations, and real-world trials with a self-reconfigurable robot, Panthera, across a range of dimensions (Section 6). The results from the experiments show that G-RRT* outperforms all the tested state-of-the-art asymptotically optimal planners, especially in high-dimensional problems. It often searches for faster initial solutions and finds better solutions with a faster convergence rate for problems seeking to minimize path length.
Since the approximation and search of G-RRT* are based on a single sample per iteration, further research interest lies in extending the idea of greedy informed sampling to batches of randomly generated samples, such as Batch Informed Trees (BIT*) (Gammell et al., 2020). Additionally, instead of relying solely on static information state spaces, we intend to develop an online version of the geometric path planner that also considers environmental changes in practical, real-world scenarios.
@@
Information on the OMPL implementations of G-RRT* will be publicly made available at https://github.com/mlsdpk/ompl.
8 Acknowledgement
This research is supported by the National Robotics Programme under its National Robotics Programme (NRP) BAU, Ermine III: Deployable Reconfigurable Robots, Award No. M22NBK0054 and also, it is supported by the National Robotics Programme under its National Robotics Programme (NRP) BAU, PANTHERA 2.0: Deployable Autonomous Pavement Sweeping Robot through Public Trials, Award No. M23NBK0065.

### v2

@@ VII Conclusion and Discussion
Anytime sampling-based planning algorithms have been proven to perform better in high-dimensional state spaces compared to graph-search approaches since these methods do not explicitly construct high-dimensional occupancy obstacles; instead, they rely on random sampling to asymptotically find optimal solutions. Algorithms like RRT* increasingly build dense approximations over the problem domain based on randomly sampled states and find continually improving solutions over time. This makes RRT* probabilistically complete and almost-surely asymptotically optimal [18].
@@
Although they are effective in continuous planning problems, their search is often inefficient since they tend to uniformly explore the entire state space (i.e., they are space-filling algorithms). In other words, these algorithms tend to spend time exploring unnecessary regions of the problem domain where solutions do not exist. This wasted computational effort becomes more significant in planning problems with higher dimensions of state space.
@@
Previous studies therefore attempt to focus the search solely on promising regions of the problem domain where solutions are likely to exist. Algorithms such as Informed RRT* directly focus sampling on a subset of the state space determined by the current solution path. This $L^{2}$ informed set contains the set of states that can potentially provide better solutions and often ensures faster convergence guarantees for planning scenarios aimed at minimizing path lengths. However, since this set is estimated using the cost of the current solution path, it can still result in larger sampling regions, especially when the current solution contains many twists and turns.
@@
We present an alternative version of the informed set, called the $L^{2}$ greedy informed set, which is approximated based on the exploited knowledge of the current solution path using the maximum admissible heuristic cost (Section III). Instead of using the cost of the current solution as an admissible heuristic for the informed set, the greedy informed set only considers the maximum heuristic cost along the current solution path. This results in smaller informed subsets and is shown to have better convergence properties than the informed set, typically in problems with complex initial solution paths.
This paper also studies the formal guarantees of the greedy informed set, analyzing how the use of heuristics affects optimality. If the true optimal solution path lies within the same homotopy class as the current solution path, the greedy informed set is guaranteed to include the states along the optimal path (Theorem 1). However, in certain planning problems where the optimal path and current solution path lie in different homotopy regions, the greedy informed set results in suboptimal solutions, failing to include the states along the optimal path (Theorem 2). This analysis shows that, depending on the nature of the problem domain, an appropriate balance between exploiting the greedy informed set and exploring with the larger informed set is essential to preserve the asymptotic optimality of informed sampling-based planners.
@@
While direct informed sampling techniques prove effective compared to traditional space-filling approaches, their effectiveness only becomes apparent once an initial solution is found. Consequently, the benefits of focused sampling diminish in complex planning problems where finding an initial solution is challenging, especially in high-dimensional state spaces. Hence, finding a fast initial solution is crucial for sampling-based planning algorithms to leverage all the benefits of informed sets. This paper also presents an algorithm that uses bidirectional search to accelerate the process of finding initial solutions and leverages the greedy informed set to focus the search using exploited knowledge of the current solution.
@@
To this end, this paper presents an almost-surely asymptotically optimal path planning algorithm that utilizes greedy heuristic techniques to address challenges in high-dimensional planning domains (Section IV). G-RRT* maintains two rapidly growing random geometric graphs (RGGs) to explore the problem state space and uses a greedy connection heuristic to guide the trees toward each other. This helps G-RRT* find faster initial solutions and often escape local minima with biased Voronoi regions.
@@
G-RRT* also focuses the search on the $L^{2}$ greedy informed set once the initial solution is found, i.e., forward and reverse search trees meet each other. This allows G-RRT* to search efficiently on the smaller subset of the problem domain with exploited knowledge of the current solution path. Moreover, a greedy biasing ratio is introduced to balance uniform sampling from the problem domain (i.e., exploration with $L^{2}$ informed set) for optimal guarantees and path-biased sampling (i.e., exploitation with $L^{2}$ greedy informed set) for accelerating convergence. As a result, G-RRT* is probabilistically complete and almost-surely asymptotically optimal (Section V). We also analyze how the greedy biasing ratio affects the worst-case performance of sampling from the greedy informed set compared to sampling only from the informed set (Theorem 3).
@@
The benefits of the greedy heuristic techniques are demonstrated on abstract planning problems, manipulation problems in simulations, and real-world trials with a self-reconfigurable robot, Panthera, across a range of dimensions (Section VI). The results from the experiments show that G-RRT* outperforms all the tested state-of-the-art asymptotically optimal planners, especially in high-dimensional problems. It often searches for faster initial solutions and finds better solutions with a faster convergence rate for problems seeking to minimize path length. We found that $\epsilon=0.9$ is often sufficient to significantly improve the convergence rate of the algorithms in most planning problems (Section VI-C); however, this value can be adjusted according to specific needs.
Since the approximation and search in G-RRT* are based on a single sample per iteration, further research interest lies in extending the ideas of greedy informed sampling to batches of randomly generated samples, such as in Batch Informed Trees (BIT*) [44]. Additionally, we are currently investigating ways to define the promising regions of the problem domain, as we believe that exploiting these regions is crucial for improving the sampling efficiency and convergence rate of sampling-based motion planners in high-dimensional planning problems.
@@
Information on the OMPL implementations of G-RRT* is publicly made available at https://github.com/mlsdpk/ompl.
Acknowledgements
The authors would like to thank Yunfan Lu for dedicating time to help with the proofs.